# FedMLLM: Federated Fine-tuning MLLM on Multimodal Heterogeneity Data

### v1

@@ 5.1 Experimental Settings
Datasets. To stay updated on the latest trends, modern society is flooded daily with social media information from various regions and across diverse scopes. However, effectively identifying and reducing misinformation and harmful content within these platforms has become a critical task in current research [9]. Moreover, social media information from different regions and privacy levels has driven research on FL in this field. Therefore, in evaluating the FedMLLM framework, we utilize two multimodal social media datasets, Hateful-Memes [17] and CrisisMMD [1], containing image-text pairs.
Configurations. For the model configuration, the efficient MiniCPM-V-2_6-int4 [37] is adopted, with Qwen2 [34] as the LLM, and LoRA parameters are set to a rank of 8, alpha of 8, and dropout of 0.05. The maximum sequence length for the model is 1200. Formatting instructions can be found in the Supplementary Material. In the FL settings, there are 50 training rounds with 10 clients, sampling 2 clients per round, each training for one epoch. The initial learning rate is set to 2e-5 and adjusted using a cosine scheduler with a warmup ratio of 0.01. In the local baseline, the number of training epochs is 5. The data split method is described in Section Sec. 3. In the non-aligned modal scenario, partitioning first uses from the aligned modal scenario, followed by the specific non-aligned modal partition method. All experiments use a single NVIDIA A40 or A100 GPU with 617M trainable parameters out of 8716M total, using around 30GB of GPU memory.
@@ 5.2 Performance Evaluation
Comparison of FL Baselines and Local. Tab. 2-Tab. 5 sequentially present the results of five comparison baselines across two downstream datasets under 12 different levels of heterogeneity in the aligned modal, missing modal, cross modal, and hybrid modal scenarios. Compared to the “Local” baseline that averages results from clients independently training on local data, the four FL baselines perform better in heterogeneous conditions, especially FedYogi [30], FedAdam [30], and FedAdagrad [30] that utilize adaptive optimizers. This indicates that collaborative training in FL, which indirectly expands the range of training data, further improves MLLM performance in heterogeneous scenarios.
Non-Aligned vs. Aligned Modal Scenarios. Fig. 5 illustrates how the performance of various FL baselines shifts from the aligned modal scenario () to non-aligned modal scenarios. In the multi-class CrisisMMD dataset, the performance of FedYogi, FedAdam, and FedAdagrad is generally lower than that in aligned modal scenarios. For example, the performance of FedYogi decreases by 6.33% when heterogeneity is compared to . However, in the binary-class Hateful-Memes dataset, FedYogi and FedAdam even show performance improvements under missing and cross modal scenarios. For instance, FedAdam improves by 3.04% when heterogeneity is I-5:T-5 compared to . On the contrary, FedAvgM [10] is more affected by data heterogeneity at than at , as shown in Tab. 2. Overall, in binary-class classification tasks, moderate modal heterogeneity functions similarly to data augmentation, enhancing model robustness to some extent and thereby improving performance. However, in multi-class classification tasks, modal heterogeneity primarily has a negative impact, leading to a decline in the performance of FL methods.
@@
Results on Modality-agnostic Strategies. Tab. 6 and Tab. 7 demonstrate the performance improvement of two modality-agnostic strategies across various non-aligned modal scenarios and FL methods, respectively. The listed non-aligned scenarios are those in which performance is most negatively affected, while the addition of both modality-agnostic strategies improves performance in these scenarios. For different FL methods, both strategies also help mitigate the negative effects of modal heterogeneity. Specifically, FedAdagrad, which is solely negatively affected by modal heterogeneity, improves with both strategies, showing a 5.01% gain on the CrisisMMD dataset in the Hybrid modal scenario. Furthermore, Tab. 8 presents ablation studies showing that prompt and regularization strategies effectively mitigate modality heterogeneity, yielding the best performance when applied together. Fig. 7 shows the ablation study for the and layers in adaptive regularization, with final values of and set to 0.1 and 4, respectively.
@@
| Prompt | Reg. | Hateful-Memes | CrisisMMD |
| 69.93 | 53.45 | ||
| ✓ | 73.51(+3.58) | 53.86(+0.41) | |
| ✓ | 73.41(+3.48) | 53.82(+0.37) | |
@@
6 Conclusion and Discussion
Conclusion. In this work, we presented a multimodal heterogeneity benchmark to evaluate the performance of federated fine-tuning for MLLMs across five FL baselines on two downstream datasets, incorporating more than ten multimodal heterogeneities across four multimodal scenarios. To tackle the challenges posed by complex multimodal heterogeneities, we have also designed two simple yet effective modality-agnostic strategies. In the proposed FedMLLM framework, we conduct a series of experiments to highlight the advantages of FL over local training, examine the impact of multimodal heterogeneities, and evaluate the effectiveness of the two modality-agnostic strategies.
Discussion. FedMLLM currently focuses on the most fundamental and essential image-text multimodal data. However, other modalities—such as video and audio—still require further exploration in future work. When introducing video modalities, especially long videos, a key challenge is achieving efficient training and deployment given the limited computational resources on client devices. Additionally, the methods used in this benchmark to construct different modality scenarios are relatively simple. We hope that future benchmark studies will incorporate more realistic data distributions to better reflect real-world scenarios.
@@ 6 Conclusion and Discussion
More Visualization of Heterogeneity. In the main text, we show the visualizations of the heterogeneity of aligned modal scenarios for the CrisisMMD dataset and the multimodal heterogeneity visualizations for the four multimodal scenarios in the Hateful-Memes dataset, as shown in Fig. 3 and Fig. 4. Furthermore, Fig. 10 and Fig. 11 supplement the visualizations of aligned modal heterogeneity and provide additional multimodal heterogeneity visualizations for the four multimodal scenarios in the Hateful-Memes dataset. Since the multimodal heterogeneity in the CrisisMMD dataset is similar to that in the Hateful-Memes dataset, redundant visualizations are omitted here for brevity.
Complete Formatting Instructions. To prepare the fine-tuning data, the data format must be adjusted according to the prompt template shown in Tab. 9. Each sample should be structured as a dictionary containing the keys: “id”, “image”, and “conversations”. The “image” key specifies the path to the image, while “conversations” includes two dictionaries: one representing the user’s instruction, which contains the question and related information, and the other representing the assistant’s (MLLM) response. Since the task primarily involves classification, the response consists of the option letter and the corresponding category. For more details on data construction, please refer to the provided code.

### v2

@@ 4.1 Experimental Settings
Baselines. In different multimodal scenarios, to evaluate the performance of various FL baseline, we refer to [48, 49] and provide four classic FL methods as baselines. FedYogi, FedAdam, and FedAdagrad [39] apply the centralized Yogi [52], Adam [21], and Adagrad [9] optimizers to the server-side aggregation in the FL framework, while FedAvgm [12] uses momentum updates on the server, similar to momentum applied on top of SGD. FedAvg [36] and FedProx [27] are not included in the benchmark due to their poor convergence, as shown in Suppl. Additionally, we consider “Zero-shot” and the local baseline for comparison. “Zero-shot” refers to the original MLLM’s test results without federated fine-tuning, while the “Local” baseline is the average test results of models trained independently on local data.
Datasets. To comprehensively evaluate different multimodal scenarios, we conduct multimodal training on four datasets from the domains of social media, natural disasters, and medical knowledge—Hateful-Memes [20], CrisisMMD [2], MedAlpaca [11], and VQA-RAD [23]. The evaluation is performed on the test sets of Hateful-Memes, CrisisMMD, VQA-RAD, and SLAKE [30], where SLAKE serves as an unseen dataset to assess the zero-shot performance of FedMLLM. The evaluation metrics include AUC, F1, and accuracy for open-ended, closed-ended, and overall answers, with accuracy being assessed by GPT-4 [1]. The dataset information is shown in Tab. 1. More details in Suppl.
Configurations. To adapt to resource-constrained FL scenarios, we chose the existing lightweight MLLM (e.g., MiniCPM-V-2_6-int4 [47]) for fine-tuning via LoRA. The LoRA parameters are set to a rank of 8, alpha of 8, and dropout of 0.05. The maximum sequence length for the model is 1200. Formatting instructions can be found in Suppl. In the FL setting, the proposed FedMLLM focuses on the cross-silo scenario [14], where, by default, 2 clients are randomly selected per round from total 10 clients. We also experimented with different numbers of clients. In the training configuration, all datasets start with an initial learning rate of 2e-5 and are updated using a cosine scheduler with a 0.01 warmup ratio over 50 rounds. Specifically, the Hateful-Memes dataset ends training after 25 rounds, the CrisisMMD dataset after 40 rounds, and the MedAlpaca and VQA-RAD datasets after 50 rounds. In the benchmark, the results are obtained by splitting with different random seeds and averaging over three runs, reporting the mean and standard deviation. The construction methods of different datasets in various multimodal scenarios are detailed in Sec. 2. More training details can be found in Suppl.
| MLLMs | Method | Comm. cost | Selected/Total Clients | |||
| 2/10 | 4/10 | 5/20 | 10/30 | |||
| MiniCPM-V-2_6int4 | Zero-shot | 0M | 66.570.00 | 66.570.00 | 66.570.00 | 66.570.00 |
| Local | 0M | 66.800.11 | 66.800.11 | 64.460.13 | 63.500.08 | |
| FL | 617M | 75.581.72 | 76.420.34 | 69.700.45 | 66.910.62 | |
| MiniCPM-Llama3-V-2_5int4 | Zero-shot | 0M | 69.290.00 | 69.290.00 | 69.290.00 | 69.290.00 |
| Local | 0M | 67.020.09 | 67.020.09 | 63.210.05 | 63.260.10 | |
| FL | 621M | 72.440.51 | 70.190.64 | 68.820.99 | 71.160.72 | |
@@ 4.2 Performance Evaluation
Comparison of different baselines. Tab. 2-Tab. 6 sequentially present the results of the “Zero-shot”, “Local”, and FL baselines on Hateful-Memes, CrisisMMD, VQA-RAD, and SLAKE across different levels of modality heterogeneity. In most scenarios, the performance of “Local” and the majority of FL baselines surpasses that of “Zero-Shot”, indicating that existing MLLMs still require further fine-tuning on private datasets for better real-world deployment. Compared to “Local”, most FL baselines show significant improvement across different scenarios, especially on the multi-class CrisisMMD dataset, where FedAvgM does not perform well in situations with strong modality heterogeneity. In the medical visual question answering task, which includes both open- and closed-ended questions, the improvement of “Local” and FL over “Zero-shot” is more pronounced in closed-ended questions. Compared to VQA-RAD, the open-ended questions in the unseen SLAKE dataset pose more challenges for “Local” and “FL”.
MLLMs, clients number, and comm. cost. As shown in Tab. 7, we experimented with lightweight MLLMs “MiniCPM-V-2_6int4” and “MiniCPM-Llama3-V-2_5int” across different client counts. For different MLLMs, the FL baseline still show a clear advantage compared to the “Zero-shot” and “Local” baselines. This federated fine-tuning approach ensures data privacy while expanding the data range beyond local data, further boosting the performance of MLLMs. As the number of clients increases, sample size per client decreases, reducing “Local” performance. However, with 30 clients, the FL baseline with “V-2_5” MLLM improves by 1.87 over “Zero-shot”, and with 20 clients, “V-2_6” improves by 3.13. This shows the FL baseline can handle client fluctuations in a cross-silo setting, with a affordable communication cost of around 620M.
@@
Results on modality-agnostic strategies. Tab. 8 and Tab. 9 demonstrate the performance improvement of two modality-agnostic strategies across various non-aligned modal scenarios and FL methods, respectively. The listed non-aligned scenarios are those in which performance is most negatively affected, while the addition of both modality-agnostic strategies improves performance in these scenarios. For different FL methods, both strategies also help mitigate the negative effects of modal heterogeneity. Specifically, FedAdagrad, which is significantly negatively affected by modality heterogeneity, improves with both strategies, showing a 5.01% gain on the CrisisMMD dataset in the Hybrid modal scenario.
Furthermore, Tab. 10 presents ablation studies showing that prompt and regularization strategies effectively mitigate modality heterogeneity, yielding the best performance when applied together. Compared to FedProx [27], which applies fixed regularization across all layers, our approach dynamically adjusts regularization strength based on modality heterogeneity, achieving superior performance in complex scenarios. Fig. 7 shows the ablation study for the and layers in improved regularization, with final values of and set to 0.1 and 4, respectively.
@@
| Prompt | Reg. | FedProx | Hateful-Memes | CrisisMMD |
| 69.93 | 53.45 | |||
| ✓ | 73.51(+3.58) | 53.86(+0.41) | ||
| ✓ | 73.41(+3.48) | 53.82(+0.37) | ||
| ✓ | ✓ | 72.31(+2.38) | 55.68(+2.23) | |
@@
5 Discussion
Number of modality. FedMLLM currently focuses on the most fundamental and essential image-text multimodal data. However, other modalities—such as video and audio—still require further exploration in future work. When introducing video modalities, especially long videos, a key challenge is achieving efficient training and deployment given the limited computational resources on clients.
Splitting of data. The data partitioning methods for the four modality scenarios are widely used and fundamental, offering researchers a simple and efficient environment for data construction and testing. However, real-world environments are much more complex, and we look forward to more benchmark studies in the future that provide more diverse and challenging multimodal scenarios.
Number of client. We focus on the cross-silo FL configuration, typically involving small institutions like hospitals with fewer than 100 clients. In our experiments, we evaluated the impact of client numbers on FL performance using 10, 20, and 30 clients, with different clients randomly selected in each round. Cross-device FL, involving thousands of resource-constrained devices like smartphones, presents significant challenges for federated fine-tuning of MLLMs and may be a future research direction.
Diversity of data. Our benchmark includes five datasets from social media, natural disasters, and medical knowledge domains, covering classification and VQA tasks. However, FedMLLM is not limited to these datasets. With the scene construction descriptions, evaluation details, and data templates provided in the main text and supplementary materials, researchers can adapt it to their own datasets for training and testing.
Adequacy of assessment. In this benchmark, we evaluated four scenarios, including zero-shot performance of the untuned MLLM, comparisons with the Local baseline and four classic FL methods, and AUC and F1 metrics for classification tasks. Two lightweight MLLMs were tested under different client configurations. For the Med-VQA task, accuracy on open-end and closed-end questions was assessed by GPT-4. Additionally, the unseen SLAKE dataset was evaluated, with open-end questions presenting a particular challenge. Future research in this area is anticipated.
Efficiency analysis. Our method enables the adaptation of general MLLMs to downstream tasks, such as hateful meme classification in social media, rescue-type classification in natural disasters, and medical visual question answering based on radiology images and questions, via federated learning with an affordable communication cost. Optimizing efficiency will be a focus of future work.
6 Conclusion
In this work, we presented a multimodal heterogeneity benchmark to evaluate the performance of federated fine-tuning for MLLMs across six FL baselines on five downstream datasets, incorporating more than ten multimodal heterogeneities across four multimodal scenarios. To tackle the challenges posed by complex multimodal heterogeneities, we have also designed two simple yet effective modality-agnostic strategies. In the proposed FedMLLM framework, we conduct a series of experiments to highlight the advantages of FL over local training, examine the impact of multimodal heterogeneities, and evaluate the effectiveness of the two modality-agnostic strategies.
@@ 6 Conclusion
Due to space limitations, this supplementary material provides further details and clarifications, as follows,
1.
Related Work
a)
Multimodal Large Language Models.
b)
Multimodal Federated Learning.
c)
Federated Learning.
a)
2.
Dataset in details.
3.
More visualizations of multimodal scenarios.
4.
More training details.
a)
Configurations.
b)
Gradient Norm and Loss Curves.
c)
Complete Formatting Instructions.
a)
1 Related Work
Multimodal Large Language Models. The rapid development of Large Language Models (LLMs) has significantly driven the research surge in Multimodal Large Language Models (MLLMs) [50]. Notably, the emergence of commercial models such as GPT-4 [37] and Gemini [41] has attracted widespread attention. Simultaneously, increasingly powerful open-source models [25, 8, 55] have also emerged. BLIP-2 [25] is well-known for its Q-former module, with promising VQA results. Building on BLIP-2 and Vicuna [8], MiniGPT-4 [55], derived from GPT-4 [37], is proposed for image-text pairs. Unlike previous methods, LLaVA [32] is trained on multimodal instruction data generated by GPT-4 to enhance its zero-shot capability.
Multimodal Federated Learning. Unlike the more established single-modal FL, Multimodal Federated Learning (MFL) has significant potential for exploration [10]. Existing MFL methods can also be categorized based on the different modal scenarios in our benchmark. In aligned modal scenarios, MFL techniques resemble those of single-modal FL, focusing on client-specific and shared knowledge [5]. Some methods also draw from multimodal learning, especially in feature fusion [10]. In missing modal scenarios, some MFL methods rely on pre-trained generative models to fill the gaps in modality [4]. In cross modal scenarios, some MFL methods employ a two-branch structure to separately constrain the learning of specific and shared modalities [46, 45, 38], while hybrid modal MFL methods utilize various graph techniques to capture complex correlations among clients [6, 7].
Federated Learning. As a classic framework for privacy protection, Federated Learning (FL) typically allows multiple parties to co-train while keeping the data local [15]. The original FedAvg [36] method aggregates model weights from different clients solely based on sample proportions, which makes it challenging to effectively address complex data heterogeneity. As a result, various variants have emerged: some constrain the update direction of the local model on the client side [27, 19], while others optimize the aggregation method on the server side [39, 12, 40].
2 Dataset
MedAlpaca [11] contains approximately 34k Q&A pairs derived from the front and back of medical flashcards, with images removed to retain only concise and targeted questions and answers. This dataset covers areas such as anatomy, physiology, pathology, and pharmacology, aiming to support the learning and understanding of basic medical sciences, clinical knowledge, and clinical skills.
VQA-RAD [23] is a dataset for Medical Visual Question Answering (Med-VQA), consisting of 2,248 Q&A pairs and 315 radiology images. The training set includes 1,797 Q&A pairs, while the test set contains 451 Q&A pairs. This Med-VQA dataset includes 11 question types: modality, abnormality, and more. The evaluation metric follows the GPT-4 [1] accuracy evaluation for open-ended and closed-ended answers, as mentioned in [31].
SLAKE [30] is an English-Chinese bilingual dataset for the Med-VQA task, containing 14k Q&A pairs and 642 images with 2,070 test pairs. The questions in this Med-VQA dataset include vision-only and knowledge-based types. According to [31], the evaluation metrics include the accuracy of open-ended and closed-end answers, as well as overall accuracy, assessed by GPT-4 [1].
3 Visualization
Modality heterogeneity. In the main text, we show the visualizations of the heterogeneity of aligned modal scenarios for the CrisisMMD dataset and the multimodal heterogeneity visualizations for the four multimodal scenarios in the Hateful-Memes dataset, as shown in Fig. 8 and Fig. 3. Furthermore, Fig. 10 and Fig. 11 supplement the visualizations of aligned modal heterogeneity and provide additional multimodal heterogeneity visualizations for the four multimodal scenarios in the Hateful-Memes dataset. Since the multimodal heterogeneity in the CrisisMMD dataset is similar to that in the Hateful-Memes dataset, redundant visualizations are omitted here for brevity.
In addition to the category heterogeneity in Hateful-Memes and CrisisMMD, we also visualize the heterogeneity brought by different category types and question types in the datasets. As shown in Fig. 9, CrisisMMD and Hateful-Memes each contain 8 rescue categories and a binary classification for malicious memes; VQA-RAD and SLAKE are two Med-VQA datasets, with 11 and 10 question categories, respectively, and there are differences in the quantities. Additionally, these two Med-VQA datasets include both open-ended and closed-ended answer types, with the quantity differences shown in Fig. 12.
4 Training Details
Configurations. The local baseline uses 5 training epochs for optimal performance, as shown in Tab. 11. The FL baselines train for 50 rounds, with 1 epoch per round. The difference in the number of epochs is normal [48, 49], as the local baseline aims to achieve the best result, while training too many epochs per round in the FL baselines may introduce more bias.
| Ep | 1 | 2 | 3 | 4 | 5 | 6 |
| C | 25.22 | 26.61 | 32.85 | 31.44 | 33.34 | 32.06 |
| H | 65.74 | 65.76 | 65.86 | 66.05 | 66.83 | 66.44 |
We train on a lightweight MLLM using LoRA, with LoRA parameters transmitted between the client and server, which is acceptable in FL. In the benchmark, the FL baselines differ only in the aggregation computation on the server side, with minimal impact on efficiency. As shown in Tab. 12, we provide the communication cost, computational cost, and total training time for Hateful-Memes (H) and CrisisMMD (C).
| Trainable / Total | Comm. cost | Train time | |
| 617M / 8716M | 617M | H: 11h 05m | C: 21h 04m |
In cross modal scenario, suppose is set to . When the missing rate is in missing modal scenario, is linearly adjusted to to adaptively adjust the strength of agreement with the global model. The hybrid modal scenario combines aligned, missing, and cross modal scenarios. Thus, the setting follows these scenarios, with no regularization in the aligned case. In practice, is calculated from the client’s private dataset.
@@ 4 Training Details
Complete Formatting Instructions. To prepare the fine-tuning data, the data format must be adjusted according to the prompt template shown in Tab. 13. Each sample should be structured as a dictionary containing the keys: “id”, “image”, and “conversations”. The “image” key specifies the path to the image, while “conversations” includes two dictionaries: one representing the user’s instruction, which contains the question and related information, and the other representing the assistant’s (MLLM) response. Since the task primarily involves classification, the response consists of the option letter and the corresponding category. For more details on data construction, please refer to the code provided in the supplementary materials.